OpenStudy (monkey*):

Which expressions are equivalent to 8•8•8•8? 8^3•8^1 1/8^-4 None of the above

rebeccaxhawaii (rebeccaxhawaii):

\[8*8*8*8=8^4\] remember \[x^a+x^b=x ^{a+b}\]